Let random variable x represent the number of heads when a fair coin is tossed two times.

a) construct a table describing probability distribution

b) determine the mean and standard deviation of x (round to 2 decimal places)
Answer #1

Solution:-

a) The table describing probability distribution is

C) 1 2 0.25 0.5 0.25 1 C) 0.5 0.5 1 C) 0.5 1 1.5 Sum

b) The mean and standard deviation of x is 1 and 1.118.

\mu = \sum x\ast P(x)

\mu = 1

\sigma =\sqrt{\left (\sum x^{2}\times P(x)} \right - \mu^{2})

\sigma =1.118